Jane’s flight from Nairobi to Guangzhou can be approximated by a 8302km vector due west. The jet that Jane is flying on has a top speed of 700km/h and there is 80km/h wind blowing in the direction of 60o south of east. What is the shortest amount of time Jane’s flight can take?
1
Expert's answer
2019-10-24T09:28:22-0400

Opposing speed of wind"=80cos\theta=40\frac{Km}{hr}"

Net speed of flight"=660\frac{Km}{hr}"

Hence, time taken"=\frac{8302}{660}=12.578hr"